# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 2.3.1 Geometry of twisted bilayer graphene
Figure 6: Commensurate TBG with rotation angle $\theta$ between the two layers. (a) Sketch of moiré pattern for (4,5) TBG, i.e., $\theta=7.34^{\circ}$. The black line shows the moiré cell containing $N=244$ C atoms. (b) K and K’ points of the (4,5) TBG reciprocal lattice. The black line shows the first Brillouin zone of the TBG, the green and magenta lines are the first Brillouin zone of the two graphene layers. (c) Zoom on the TBG Brillouin zone. (d) Intraband velocity at Dirac energy $E_{D}$ (computed from Eq. (73) at the K point) versus the rotation angle $\theta$, for two TB models corresponding to $V_{MLG}=1.09\cdot 10^{6}$ ms-1 which correctly reproduces the position of the Van Hove singularity found experimentally in TBG as a function of the angle [97], and $V_{MLG}=0.79\cdot 10^{6}$ ms-1, which reproduces DFT calculations [90, 94], respectively. Some data from Refs. [90, 94, 106]. (e,f,g) Electronic structure of (12,13) TBG ($\theta=2.66^{\circ}$): SK-TB local DOS (LDOS) (arbitrary units) in real space for energy close to $E_{D}$, (e) SK-TB calculation (f) analytical model (Eq. (25)) [125]; and (g-left) SK-TB Bands, the line color is the average layer participation ratio $p^{*}$, $p^{*}=0.5/(P_{1}^{2}+P_{2}^{2})$, where $P_{i}$ it the weight of the eigenstate on layer $i$ [29]. $p^{*}=0.5$ when eigenstates are located on one layer only, and $p^{*}=1$ when eigenstates are equally distributed on the two layers. (g-right) SK-TB total DOS, LDOS at the center of an AA stacking region, and (blue line) integrated total DOS with respect to the neutral situation (half filling). (e,g) are calculated with the SK-TB model corresponding to a monolayer intraband velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation.
A single layer of graphene consists of carbon atoms arranged in a honeycomb lattice such that the unit cell includes two carbon atoms. Our starting point is AA bilayer graphene, where all the atoms in one layer are on top of an atom in the other layer; we choose the rotation origin O at an atomic site, and one of the two layers is rotated by an angle $\theta$ with respect to the other layer. A periodic commensurate bilayer structure parameterized by two integers $n$, $m$ can be constructed using the method of Refs. [89, 98, 99, 90, 100]. The rotation angle for such a commensurate structure is then given by,
@@
and the vectors of the TBG superlattice are $\mathbf{t}_{1}=n\mathbf{a}_{1}+m\mathbf{a}_{2}$ and $\mathbf{t}_{2}=-m\mathbf{a}_{1}+(m+n)\mathbf{a}_{2}$, where $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the lattice vectors of the non-rotated single graphene layer, $\mathbf{a}_{1,2}=a(\sqrt{3},\pm 1)/2$, with $a=2.456$. The number of atoms in the moiré cell is given by $N=4(n^{2}+m^{2}+mn)$. The rotation angle $\theta$ is a good parameter to describe the system, but the number of atoms $N$ is not since cells of equivalent size can be found for different angles. For $\theta$ values less than $\sim 15^{\circ}$, TBG forms a moiré pattern with (pseudo)period,
$P=\frac{a}{2\sqrt{3}\sin(\theta/2)}.$ (18)
When $|n-m|=1$, the periodic unit cell contains 1 moiré cell, and the smallest commensurable structure with rotation angle close to a given small $\theta$ is obtained for TBG $(n,m=n+1)$ [100]. Figure 6(a) shows the resulting moiré pattern for $(n,m)=(4,5)$. AA stacking regions are at the corners of the moiré cell. Along its long diagonal we find Bernal AB and BA stacking regions, where half of the atoms in one layer are on top of an atom in the other layer.
It is also possible to add translation to the rotated layer in addition to rotation. For intermediate angles, this can slightly affect the electronic structure without qualitatively changing the results. For small angles, when the moiré cell is very large, such a translation does not significantly modify the electronic structure, in line with what the crystallographic analysis suggests [100, 101].
For $\theta$ angle values that do not correspond to a commensurate structure, e.g. $\theta=20^{\circ}$, $30^{\circ}$ (…), the TBGs are quasiperiodic structures [210].
@@ 2.3.2 Tight-binding electronic structure of twisted bilayer graphene
Given the large number of atoms per TBG cell, DFT studies of the electronic structure were not originally possible for small angles $\theta$. The first theoretical studies were therefore carried out using simplified models. Two main approaches have been used since the beginning of research on TBGs. One approach is based on tight-binding models that include $p_{z}$ orbitals of all carbon atoms [89, 90, 92, 91, 93, 94, 95]. Following this approach, we proposed in 2010 a Slater-Koster tight binding (SK-TB) Hamiltonian [90, 94] which has been reused in many works in the literature (see for examples Refs. [211, 212, 213, 214, 124, 215, 216, 217, 218, 219, 128, 129, 133]). Other TB models have been also proposed (see, e.g., Ref. [113]). Another approach, originally developed by R. Bistritzer and A. H. MacDonald [91, 93] consists in a continuum model for the low-energy bands by coupling the Dirac cones of the two graphene layers. This continuum model has been enormously successful and has led to major advances in our understanding of the properties of flat bands. Some examples are Refs. [220, 221, 222, 223, 224, 225, 138, 226, 227, 228, 229, 230, 231, 232]. DFT studies for TBGs close to the magic angle have been carried out [107, 233, 121, 218], as well as other DFT-based TB models [234] and approaches using Wannier functions [235, 116, 236, 237]. More recently, a topological heavy-fermion model, combining correlated localized flat-band ($f$) electrons and itinerant conduction electrons, has also been proposed to understand the unconventional electronic behavior of magic-angle electron states [21, 130]. It is not possible in this review to cite and detail all these approaches, we will focus on the approach that we have used and developed based on a SK-TB Hamiltonian.
In order to be used numerically for systematic study, an SK-TB model requires a commensurable structure (Sec. 2.3.1) and angle-independent Slater-Koster parameters. The non-interacting SK-TB Hamiltonian is [90, 94],
@@
where $c_{i}^{\dagger}$ and $c_{i}$ are the creation and annihilation operators of an electron on orbital $p_{z}$ at site $i$, and $\langle i,j\rangle$ is the double sum on indices $i$ and $j$ with $i\neq j$. The coupling matrix element, $t_{ij}$, between two $p_{z}$ orbitals located at $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$ is calculated by the Slater-Koster formula [238],
@@
where $n$ is the direction cosine of $\mathbf{r}_{ij}=\mathbf{r}_{j}-\mathbf{r}_{i}$ along the $z$ axis and $r_{ij}$ the distance between the orbitals, $n={z_{ij}}/{r_{ij}}$, with $z_{ij}$ is the $z$-coordinate of $\mathbf{r}_{ij}$. $z_{ij}$ is either equal to zero or to a constant because the two graphene layers have been kept flat in our model (rigid TBGs). The same distance dependence of the two SK parameters, $V_{pp\pi}$ and $V_{pp\sigma}$, is used:
@@
where $a_{0}$ is the nearest-neighbor distance within a layer, $a_{0}=a\sqrt{3}=1.418$, and $a_{1}$ is the interlayer distance, $a_{1}=3.349$. First neighbors interaction in a plane is taken equal to the commonly used value, $V_{pp\pi}^{0}=-t_{0}=-2.7$ eV [239]. Second neighbors interaction $t_{0}^{\prime}$ in a plane is set to $0.1t_{0}$ [239] which fixes the value of the ratio $q_{\pi}/{a_{0}}$ in equation (21). The interlayer coupling between two $p_{z}$ orbitals in $\pi$ configuration is $V_{pp\sigma}^{0}$. $V_{pp\sigma}^{0}$ is fixed to obtain a good fit with DFT calculation around the Dirac energy in AA stacking and AB bernal stacking which results in $V_{pp\sigma}^{0}=0.48$ eV. We chose the same coefficient of the exponential decay for $V_{pp\pi}$ and $V_{pp\sigma}$,
@@
with $a=2.456$ the distance between second neighbors in a monolayer of graphene. In (21), a smooth cut-off function [240] is introduced to avoid the unrealistic cut-off effects that can lead to the formation of false mini-gaps [125],
@@
with $l_{c}=0.5$ a.u. $=0.265$, and the cut-off distance is $r_{c}=2.5a=6.14$. Without asymmetric doping, all $p_{z}$ orbitals have the same on-site energy $\epsilon_{i}$ (Eq. (19)). $\epsilon_{i}$ is set to $=-0.7834$ eV so that the energy $E_{D}$ of the Dirac point K is equal to zero in monolayer graphene ($E_{D(MLG)}=0$). $\epsilon_{i}$ is not zero because the intralayer coupling between atoms beyond first neighbors breaks the electron/hole symmetry and then shifts $E_{D}$.
As explained above, the SK parameters given below have been tuned to obtain good agreement with low-energy DFT band states for the monolayer graphene, Bernal AB bilayer, and the AA bilayer [90, 94]. However, it is well known that these bands yield an intraband Fermi velocity of monolayer graphene, $V_{MLG}=0.79\cdot 10^{6}$ ms-1, which does not correspond to the experimental values. Consequently, in order to make comparisons with experimental results, in particular STM spectroscopy results, we have used a SK-TB model with the same parameters as above except $V_{pp\pi}^{0}=-3.7$ eV [97], to get $V_{MLG}=1.09\cdot 10^{6}$ ms-1. That SK-TB model has been used to successfully reproduce many experimental results for various angles [97, 241, 104, 242, 106, 243]. It works very well for TBGs where the two graphene layers remain flat, with or without strain or heterostrain (Sec. 2.3.5). However it is not clear that it has been well validated by comparison with DFT calculations when out-of-plane relaxation is switched on.
For rigid TBGs, i.e.,  TBGs without atomic relaxation, theoretical studies have predicted [89, 90, 92, 91, 93, 94, 95] the existence of three domains depending on the $\theta$ values:
@@
    For large rotation angles ($\theta>20^{\circ}$) the layers are “decoupled” (in the sense that the electronic eigenstates are located in only one layer) and then behave as two isolated graphene layers.
@@
    For intermediate angles, $\sim 2^{\circ}<\theta<20^{\circ}$, the dispersion, around the Fermi energy $E_{F}$, remains linear, but the velocity is reduced (see Fig. 6(d,g)). Several analytical models [89, 93, 95, 125] have been proposed to simulate this reduction, for example a perturbation of the Dirac cone for intermediate angles leads to a velocity ratio between TBG and the monolayer, [244, 125]
@@
    where $t_{k}\approx 0.1$ eV is the interlayer coupling term between the closest Dirac cone states of the two layers. Consequently, the energies $E_{-}$ and $E_{+}$ of the two van Hove singularities are shifted to the Dirac energy $E_{D}$ when $\theta$ decreases, as it has been shown experimentally [205, 97, 241]. As $\theta$ decreases, the low-energy states are more and more localized in the AA stacking region of the moiré pattern, and thus the local DOS (LDOS) in the AA regions increases (Fig. 6(e,f,g)). The analytical model based on a perturbation of the Dirac cone gives the variation of the LDOS with respect to the position $\mathbf{r}$ in real space, [244, 125]
@@
    For the lowest $\theta$, with $\theta<\sim 2^{\circ}$ (Fig. 7), almost flat bands appear and result in electronic localization in AA stacking regions: states of similar energies, belonging to the Dirac cones of the two layers interact strongly. In this regime, the velocity of states at the Dirac point goes to almost zero for specific angles, the so-called magic angles [90, 93, 94]. The following sections are devoted to the properties of these TBGs, which are characterized by low-energy flat bands.

### v2

@@ 2.3.1 Geometry of twisted bilayer graphene
Figure 6: Commensurate TBG with rotation angle $\theta$ between the two layers. (a) Sketch of moiré pattern for (4,5) TBG, i.e., $\theta=7.34^{\circ}$. The black line shows the moiré cell containing $N=244$ C atoms. (b) K and K’ points of the (4,5) TBG reciprocal lattice. The black line shows the first Brillouin zone of the TBG, the green and magenta lines are the first Brillouin zone of the two graphene layers. (c) Zoom on the TBG Brillouin zone. (d) Intraband velocity at Dirac energy $E_{D}$ (computed from Eq. (77) at the K point) versus the rotation angle $\theta$, for two TB models corresponding to $V_{MLG}=1.09\cdot 10^{6}$ ms-1 which correctly reproduces the position of the Van Hove singularity found experimentally in TBG as a function of the angle Brihuega12 , and $V_{MLG}=0.79\cdot 10^{6}$ ms-1, which reproduces DFT calculations Trambly10 , Trambly12 , respectively. Some data from Refs. Trambly10 , Trambly12 , Mesple21 . (e,f,g) Electronic structure of (12,13) TBG ($\theta=2.66^{\circ}$): SK-TB local DOS (LDOS) (arbitrary units) in real space for energy close to $E_{D}$, (e) SK-TB calculation (f) analytical model (Eq. (25)) Omid20 ; and (g-left) SK-TB Bands, the line color is the average layer participation ratio $p^{*}$, $p^{*}=0.5/(P_{1}^{2}+P_{2}^{2})$, where $P_{i}$ it the weight of the eigenstate on layer $i$ Trambly16 . $p^{*}=0.5$ when eigenstates are located on one layer only, and $p^{*}=1$ when eigenstates are equally distributed on the two layers. (g-right) SK-TB total DOS, LDOS at the center of an AA stacking region, and (blue line) integrated total DOS with respect to the neutral situation (half filling). (e,g) are calculated with the SK-TB model corresponding to a monolayer intraband velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation.
A single layer of graphene consists of carbon atoms arranged in a honeycomb lattice such that the unit cell includes two symmetrically equivalent carbon atoms. Our starting point is AA bilayer graphene, where all the atoms in one layer are on top of an atom in the other layer; we choose the rotation origin O at an atomic site, and one of the two layers is rotated by an angle $\theta$ with respect to the other layer. A periodic commensurate bilayer structure parameterized by two integers $n$, $m$ can be constructed using the method of Refs. LopesdosSantos07 , Campanera07 , Mele10 , Trambly10 , Gratias23 . The rotation angle for such a commensurate structure is then given by,
@@
and the vectors of the TBG superlattice are $\mathbf{t}_{1}=n\mathbf{a}_{1}+m\mathbf{a}_{2}$ and $\mathbf{t}_{2}=-m\mathbf{a}_{1}+(m+n)\mathbf{a}_{2}$, where $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the lattice vectors of the non-rotated single graphene layer, $\mathbf{a}_{1,2}=a(\sqrt{3},\pm 1)/2$, with $a=2.456$ Å. The number of atoms in the moiré cell is given by $N=4(n^{2}+m^{2}+mn)$. The rotation angle $\theta$ is a good parameter to describe the system, but the number of atoms $N$ is not since cells of equivalent size can be found for different angles. For $\theta$ values less than $\sim 15^{\circ}$, TBG forms a moiré pattern with (pseudo)period,
$P=\frac{a}{2\sin(\theta/2)}.$ (18)
When |nm|=1|n-m|=1, the periodic unit cell contains 1 moiré cell, and the smallest commensurable structure with rotation angle close to a given small $\theta$ is obtained for TBG $(n,m=n+1)$ Gratias23 . Figure 6(a) shows the resulting moiré pattern for $(n,m)=(4,5)$. AA stacking regions are at the corners of the moiré cell. Along its long diagonal we find Bernal AB and BA stacking regions, where half of the atoms in one layer are on top of an atom in the other layer.
It is also possible to add translation to the rotated layer in addition to rotation. For intermediate angles, this can slightly affect the electronic structure without qualitatively changing the results. For small angles, when the moiré cell is very large, such a translation does not significantly modify the electronic structure, in line with what the crystallographic analysis suggests Gratias23 , Gratias20 .
For $\theta$ angle values that do not correspond to a commensurate structure, e.g. $\theta=20^{\circ}$, $30^{\circ}$ (…), the TBGs are quasiperiodic structures Gratias24 .
@@ 2.3.2 Tight-binding electronic structure of twisted bilayer graphene
Given the large number of atoms per TBG cell, DFT studies of the electronic structure were not originally possible for small angles $\theta$. The first theoretical studies were therefore carried out using simplified models. Two main approaches have been used since the beginning of research on TBGs. One approach is based on tight-binding models that include $p_{z}$ orbitals of all carbon atoms LopesdosSantos07 , Trambly10 , SuarezMorell10 , Bistritzer10 , Bistritzer11 , Trambly12 , Santos12 . Following this approach, we proposed in 2010 a Slater-Koster tight binding (SK-TB) Hamiltonian Trambly10 , Trambly12 which has been reused in many works in the literature (see for examples Refs. Moon12 , Moon13 , Wang12 , Gargiulo17 , Andelkovic18 , Kerelsky2019 , Klebl19 , Goodwin19 , Pathak22 , Nguyen22 , Sinha24 , guerrero24_tBLG_disorder_Kubo , Guerrero25 ). Other TB models have been also proposed (see, e.g., Ref. Lin18 ). Another approach, originally developed by R. Bistritzer and A. H. MacDonald Bistritzer10 , Bistritzer11 , consists in a continuum model for the low-energy bands by coupling the Dirac cones of the two graphene layers. This continuum model has been enormously successful and has led to major advances in our understanding of the properties of flat bands. Some examples are Refs. Weckbecker16 , Guinea18 , Tarnopolsky19 , Guinea19 , tbgI , Mora19 , andrei2020graphene , XIONG2023129048 , Pantaleon21 , Mannai21 , Haddad23 , Cances23 , Escudero24 , Yu2025 . DFT studies for TBGs close to the magic angle have been carried out Uchida14 , Yndurain19 , Leconte22 , Pathak22 , as well as other DFT-based TB models Jung14 and approaches using Wannier functions Fang16 , Koshino18 , Kang18 , Vafek19 . More recently, a topological heavy-fermion model, combining correlated localized flat-band ($f$) electrons and itinerant conduction electrons, has also been proposed to understand the unconventional electronic behavior of magic-angle electron states Song22 , Rai24 . It is not possible in this review to cite and detail all these approaches, we will focus on the approach that we have used and developed based on a SK-TB Hamiltonian.
In order to be used numerically for systematic study, an SK-TB model requires a commensurable structure (Sec. 2.3.1) and angle-independent Slater-Koster parameters. The non-interacting SK-TB Hamiltonian is Trambly10 , Trambly12 ,
@@
where $c_{i}^{\dagger}$ and $c_{i}$ are the creation and annihilation operators of an electron on orbital $p_{z}$ at site $i$, and $\langle i,j\rangle$ is the double sum on indices $i$ and $j$ with $i\neq j$. The coupling matrix element, $t_{ij}$, between two $p_{z}$ orbitals located at $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$ is calculated by the Slater-Koster formula Slater54 ,
@@
where $n$ is the direction cosine of $\mathbf{r}_{ij}=\mathbf{r}_{j}-\mathbf{r}_{i}$ along the $z$ axis and $r_{ij}$ the distance between the orbitals, $n={z_{ij}}/{r_{ij}}$, with $z_{ij}$ is the $z$-coordinate of $\mathbf{r}_{ij}$. $z_{ij}$ is either equal to zero or to a constant because the two graphene layers have been kept flat in our model. The same distance dependence of the two SK parameters, $V_{pp\pi}$ and $V_{pp\sigma}$, is used:
@@
where $a_{0}$ is the nearest-neighbor distance within a layer, $a_{0}=a\sqrt{3}=1.418$ Å, and $a_{1}$ is the interlayer distance, $a_{1}=3.349$ Å. First neighbors interaction in a plane is taken equal to the commonly used value, $V_{pp\pi}^{0}=-t_{0}=-2.7$ eV Castro09_RevModPhys . Second neighbors interaction $t_{0}^{\prime}$ in a plane is set to $0.1t_{0}$ Castro09_RevModPhys which fixes the value of the ratio $q_{\pi}/{a_{0}}$ in equation (21). The interlayer coupling between two $p_{z}$ orbitals in $\pi$ configuration is $V_{pp\sigma}^{0}$. $V_{pp\sigma}^{0}$ is fixed to obtain a good fit with DFT calculation around the Dirac energy in AA stacking and AB bernal stacking which results in $V_{pp\sigma}^{0}=0.48$ eV Trambly10 , Trambly12 . We chose the same coefficient of the exponential decay for $V_{pp\pi}$ and $V_{pp\sigma}$,
@@
with $a=2.456$ Å the distance between second neighbors in a monolayer of graphene. In (21), a smooth cut-off function Mehl96 is introduced to avoid the unrealistic cut-off effects that can lead to the formation of false mini-gaps Omid20 ,
@@
with $l_{c}=0.5$ a.u. =0.265=0.265 Å, and the cut-off distance is $r_{c}=2.5a=6.14$ Å. Without asymmetric doping, all $p_{z}$ orbitals have the same on-site energy $\epsilon_{i}$ (Eq. (19)). $\epsilon_{i}$ is set to =0.7834=-0.7834 eV so that the energy $E_{D}$ of the Dirac point K is equal to zero in monolayer graphene ($E_{D(MLG)}=0$). $\epsilon_{i}$ is not zero because the intralayer coupling between atoms beyond first neighbors breaks the electron/hole symmetry and then shifts $E_{D}$.
As explained above, the SK parameters given below have been tuned to obtain good agreement with low-energy DFT band states for the monolayer graphene, Bernal AB bilayer, and the AA bilayer Trambly10 , Trambly12 . However, it is well known that these bands yield an intraband Fermi velocity of monolayer graphene, $V_{MLG}=0.79\cdot 10^{6}$ ms-1, which does not correspond to the experimental values. Consequently, in order to make comparisons with experimental results, in particular STM spectroscopy results, we have used a SK-TB model with the same parameters as above except $V_{pp\pi}^{0}=-3.7$ eV Brihuega12 , to get $V_{MLG}=1.09\cdot 10^{6}$ ms-1. That SK-TB model has been used to successfully reproduce many experimental results for various angles Brihuega12 , Cherkez15 , Huder18 , Huder18b , Mesple21 , Mesple23 . It works very well for TBGs where the two graphene layers remain flat, with or without in-plane relaxation or heterostrain (Sec. 2.3.5).
The best proof of the validity of this SK-TB model for describing real TBGs, and in particular those at the first magic angle, is provided by the excellent agreement between experimental LDOS measurements (by STM) and TB calculations without heterostrain Brihuega12 , Cherkez15 , and in the presence of various heterostrains Huder18 , Mesple21 . These comparisons also show that, contrary to what is often stated in the literature, it does not seem necessary to introduce out-of-plane relaxation into the calculations to simulate the experimental results (for twist angles greater than or equal to the first magic angle, see also Sec. 2.3.5). We also have shown an excellent agreement between this SK-TB model and STM experiments for the effect of a strong in-plane relaxation in the case of super-moiré of several million atoms formed by biaxial heterostrain Mesple23 . It is therefore reasonable to assume that this model is valid for analyzing the effects of heterostrain and in-plane relaxation. However, it is not clear that this SK-TB model is correct in the presence of out-of-plane relaxation, as to our knowledge it has not been validated by detailed comparisons with DFT calculations. Indeed it is possible that the exponential decay (Eq. (21)) overestimates the effect of the out-of-plane relaxation. This is why we do not discuss the effect of out-of-plane relaxation in detail in this review. However, this does not seem to be a significant issue for the study of TBGs around the first magic angle, as comparisons with the experimental results cited above show that the SK-TB model without out-of-plane relaxation is valid.
For rigid TBGs, i.e., TBGs without atomic relaxation, theoretical studies have predicted LopesdosSantos07 , Trambly10 , SuarezMorell10 , Bistritzer10 , Bistritzer11 , Trambly12 , Santos12 the existence of three domains depending on the $\theta$ values:
@@
    For large rotation angles ($\theta>20^{\circ}$) the layers are “decoupled” (in the sense that the electronic eigenstates are located in only one layer) and then behave as two isolated graphene layers. This result, which now seems to surprise no one, is not so easy to understand and it has aroused interest before 2010 when it was established experimentally and by DFT calculations (see, e.g. Hass08 ). TB calculations show that to obtain this monolayer behavior, it is essential to take into account a large number of interlayer couplings, i.e., a large cut-off distance $r_{c}$ in (23). Otherwise, a non-physical minigap appears in place of the Dirac cone Trambly16 . In the monolayer, the Dirac cone comes from the triangular symmetry and the symmetry between the two sub-lattices of graphene (atoms A and B). One might think that the latter is broken in the TBGs, but it seems that long-range interlayer coupling restores it on average.
@@
    For intermediate angles, $\sim 2^{\circ}<\theta<20^{\circ}$, the dispersion, around the Fermi energy $E_{F}$, remains linear, but the velocity is reduced (see Fig. 6(d,g)). Several analytical models LopesdosSantos07 , Bistritzer11 , Santos12 , Omid20 have been proposed to simulate this reduction, for example a perturbation of the Dirac cone for intermediate angles leads to a velocity ratio between TBG and the monolayer, Omid_these , Omid20
@@
    where $t_{k}\approx 0.1$ eV is the interlayer coupling term between the closest Dirac cone states of the two layers. Consequently, the energies $E_{-}$ and $E_{+}$ of the two van Hove singularities are shifted to the Dirac energy $E_{D}$ when $\theta$ decreases, as it has been shown experimentally Luican11 , Brihuega12 , Cherkez15 . As $\theta$ decreases, the low-energy states are more and more localized in the AA stacking region of the moiré pattern, and thus the local DOS (LDOS) in the AA regions increases (Fig. 6(e,f,g)). The analytical model based on a perturbation of the Dirac cone gives the variation of the LDOS with respect to the position $\mathbf{r}$ in real space, Omid_these , Omid20
@@
    For the lowest $\theta$, with $\theta<\sim 2^{\circ}$ (Fig. 7), almost flat bands appear and result in electronic localization in AA stacking regions: states of similar energies, belonging to the Dirac cones of the two layers interact strongly. In this regime, the velocity of states at the Dirac point goes to almost zero for specific angles, the so-called magic angles Trambly10 , Bistritzer11 , Trambly12 . The following sections are devoted to the properties of these TBGs, which are characterized by low-energy flat bands.